If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?


Answer by ptaylor(2198)

Let x=amount of time it takes all three working together to mix 20 drinks
Steven mixes at the rate of 1/5 of the job ( mixing 20 drinks) per minute
Sue mixes at the rate of 1/10 of the job per minute
Jack mixes at the rate of 1/15 of the job per minute
Together they mix at the rate of 1/5 + 1/10 + 1/15 multiply each term by 30/30
6/30 + 3/30 +2/30=11/30
So, together they mix at the rate of 11/30 of the job per minute. Our equation to solve then is:
(11/30)x=1 (1 job or 20 drinks, that is)
multiply each side by 30
11x=30 divide each side by 11
x=2.727272727--- minutes

Hope this helps---ptaylor